OpenStudy (anonymous):

What is the instantaneous rate of change for tanx with respect to x at x = pi/6?

OpenStudy (anonymous):

reqd. rate of change = d(tan x) dx = sec^x when x = pi/6 rate of change = sec^2 (pi/6) = 4/3